OpenStudy (anonymous):

How would I write the standard equation for the circle center (-6, 7), r = 9?

OpenStudy (rajee_sam):

isn't it the same thing we did earlier. They have given you the center and the radius . Put it in the standard form I gave you earlier. You are smart. You should not be posting the same type of question twice.

OpenStudy (anonymous):

Oops, sorry I wrote the wrong thing! My bad.

OpenStudy (anonymous):

I have one quick question @rajee_sam. What if it was (-6,-8) and it passed through (0,0). What would my radius be?

OpenStudy (rajee_sam):

Your distance between these two points will be the radius. Do you know the distance formula?

OpenStudy (rajee_sam):



OpenStudy (rajee_sam):



OpenStudy (rajee_sam):

@Nvidia64

OpenStudy (anonymous):

=100

OpenStudy (rajee_sam):

100 is r², so r = 10

OpenStudy (anonymous):

So would it be (x+6)^2 + (y+8)^2 = 100? or (x-6)^2 + (y-8)^2 = 10?

OpenStudy (rajee_sam):

the center is -6,-8 So ( x - (-6) )² + ( y - (-8) )² = 100 which will be ( x+6)² + ( y+8)² = 100